# Superpixel Graph Contrastive Clustering with Semantic-Invariant Augmentations for Hyperspectral Images

### v1

@@ IV-B Experiment Setup
Dataset Metric K-means $FCM$ SSC SuperPCA SSSC S5C GCOT SLGSC DGAE NCSC SPGCC
IP OA .3737±.0050$\bullet$ .$3159\textpm.0035$4769±.0060$\bullet$ .5321±.0349$\bullet$ .5130±.0238$\bullet$ .4429±.0270$\bullet$ .5030±.0010$\bullet$ .5859±.0162$\bullet$ .5879±.0312$\bullet$ .5342±.0185$\bullet$ .6759±.0136
AA .4152±.0167$\bullet$ .$3541\textpm.0094$4620±.0413$\bullet$ .5099±.0724$\bullet$ .5483±.0391$\bullet$ .4876±.0300$\bullet$ .3738±.0031$\bullet$ .5703±.0254$\bullet$ .5218±.0471$\bullet$ .4382±.0294$\bullet$ .6274±.0346
Kappa .3045±.0045$\bullet$ .$2586\textpm.0035$4173±.0079$\bullet$ .4858±.0360$\bullet$ .4551±.0264$\bullet$ .3817±.0320$\bullet$ .4245±.0011$\bullet$ .5486±.0170$\bullet$ .5403±.0336$\bullet$ .4769±.0195$\bullet$ .6385±.0151
NMI .4394±.0034$\bullet$ .$4038\textpm.0013$5738±.0071$\bullet$ .6639±.0139$\bullet$ .5918±.0146$\bullet$ .5558±.0114$\bullet$ .5177±.0007$\bullet$ .6395±.0065$\bullet$ .6441±.0113$\bullet$ .5675±.0104$\bullet$ .6767±.0116
ARI .2186±.0018$\bullet$ .$1809\textpm.0013$3136±.0063$\bullet$ .4061±.0387$\bullet$ .3270±.0163$\bullet$ .3014±.0215$\bullet$ .3322±.0004$\bullet$ .4073±.0091$\bullet$ .4712±.0403$\bullet$ .3910±.0101$\bullet$ .5654±.0174
F1 .3068±.0039$\bullet$ .$2594\textpm.0012$3923±.0064$\bullet$ .4704±.0343$\bullet$ .4052±.0126$\bullet$ .3810±.0187$\bullet$ .4291±.0003$\bullet$ .4655±.0079$\bullet$ .5303±.0367$\bullet$ .4624±.0093$\bullet$ .6126±.0150
Precision .3340±.0055$\bullet$ .$3322\textpm.0013$4200±.0082$\bullet$ .5351±.0488$\bullet$ .4288±.0265$\bullet$ .4108±.0247$\bullet$ .3634±.0004$\bullet$ .5819±.0168$\bullet$ .5815±.0362$\bullet$ .4842±.0099$\bullet$ .6930±.0254
Recall .2840±.0106$\bullet$ .$2128\textpm.0011$3683±.0118$\bullet$ .4218±.0374$\bullet$ .3850±.0122$\bullet$ .3558±.0194$\bullet$ .5238±.0003$\bullet$ .3881±.0077$\bullet$ .4879±.0389$\bullet$ .4426±.0128$\bullet$ .5491±.0107
Purity .5117±.0021$\bullet$ .$4995\textpm.0033$5705±.0119$\bullet$ .6862±.0269$\bullet$ .6031±.0336$\bullet$ .5644±.0201$\bullet$ .5471±.0009$\bullet$ .6995±.0133$\bullet$ .7023±.0194$\bullet$ .6109±.0175$\bullet$ .7427±.0170
Time (s) 22.6 $31.3$ 20.3 6.4 2 1.4 146.1 108 25.5 77.5 31.1
SA OA .6697±.0013$\bullet$ .$5752\textpm.0312$7188±.0092$\bullet$ .7258±.0606$\bullet$ .7587±.0170$\bullet$ .6120±.0171$\bullet$ OOM .8209±.0228$\bullet$ .6683±.0267$\bullet$ .7647±.0254$\bullet$ .8447±.0102
AA .6856±.0055$\bullet$ .$6316\textpm.0270$5794±.0113$\bullet$ .6639±.0553$\bullet$ .6330±.0252$\bullet$ .6103±.0289$\bullet$ OOM .7125±.0071 .6577±.0430$\bullet$ .6109±.0093$\bullet$ .7216±.0146
Kappa .6328±.0010$\bullet$ .$5363\textpm.0316$6876±.0101$\bullet$ .6936±.0692$\bullet$ .7304±.0191$\bullet$ .5765±.0186$\bullet$ OOM .8011±.0241$\bullet$ .6323±.0283$\bullet$ .7374±.0272$\bullet$ .8269±.0113
NMI .7245±.0015$\bullet$ .$6935\textpm.0099$8051±.0065$\bullet$ .8267±.0143$\bullet$ .8344±.0118$\bullet$ .7623±.0097$\bullet$ OOM .8757±.0058$\circ$ .7803±.0151$\bullet$ .8361±.0205$\bullet$ .8688±.0056
ARI .5255±.0018$\bullet$ .$4507\textpm.0253$6882±.0073$\bullet$ .6639±.0647$\bullet$ .6800±.0154$\bullet$ .5014±.0135$\bullet$ OOM .7164±.0227$\bullet$ .5618±.0310$\bullet$ .6902±.0561$\bullet$ .7598±.0146
F1 .5762±.0020$\bullet$ .$5017\textpm.0249$7214±.0064$\bullet$ .7030±.0543$\bullet$ .7149±.0135$\bullet$ .5491±.0118$\bullet$ OOM .7458±.0208$\bullet$ .6069±.0314$\bullet$ .7240±.0507$\bullet$ .7850±.0130
Precision .5498±.0031$\bullet$ .$5500\textpm.0157$6903±.0090$\bullet$ .6311±.1043$\bullet$ .6683±.0170$\bullet$ .5852±.0183$\bullet$ OOM .7330±.0239$\bullet$ .5998±.0233$\bullet$ .6719±.0344$\bullet$ .7633±.0136
Recall .6053±.0086$\bullet$ .$4628\textpm.0401$7555±.0071$\bullet$ .8085±.0249 .7689±.0155$\bullet$ .5174±.0091$\bullet$ OOM .7604±.0374$\bullet$ .6227±.0868$\bullet$ .7860±.0742 .8081±.0153
Purity .6736±.0018$\bullet$ .$6707\textpm.0096$7613±.0142$\bullet$ .7430±.0523$\bullet$ .7715±.0152$\bullet$ .7068±.0166$\bullet$ OOM .8357±.0134 .7216±.0136$\bullet$ .7779±.0240$\bullet$ .8455±.0095
Time (s) 109.5 $155.2$ 11.7 92.1 12.4 2.4 OOM 47 360.2 2408.3 150.6
PU OA .5341±.0000$\bullet$ .$5162\textpm.0100$5517±.0000$\bullet$ .4666±.0177$\bullet$ .4877±.0066$\bullet$ .4601±.0058$\bullet$ OOM .5718±.0029$\bullet$ .5953±.0408$\bullet$ OOM .6268±.0048
AA .5259±.0001$\bullet$ .$\uline{.5531\textpm.0085}.$4881±.0000$\bullet$ .5205±.0228$\bullet$ .5073±.0227$\bullet$ .4606±.0084$\bullet$ OOM .5367±.0029$\bullet$ .5180±.0303$\bullet$ OOM .5903±.0053
Kappa .4336±.0000$\bullet$ .$4209\textpm.0110$4563±.0000$\bullet$ .3610±.0187$\bullet$ .4002±.0067$\bullet$ .3476±.0036$\bullet$ OOM .4822±.0031$\bullet$ .4994±.0397$\bullet$ OOM .5491±.0056
NMI .5330±.0000$\bullet$ .$5284\textpm.0081$5148±.0000$\bullet$ .4424±.0004$\bullet$ .4922±.0098$\bullet$ .4741±.0028$\bullet$ OOM .5389±.0043$\bullet$ .5348±.0208$\bullet$ OOM .5792±.0114
ARI .3213±.0000$\bullet$ .$3052\textpm.0049$3526±.0000$\bullet$ .2542±.0100$\bullet$ .2960±.0063$\bullet$ .2897±.0127$\bullet$ OOM .4276±.0025$\bullet$ .4483±.0515 OOM .4630±.0091
F1 .4543±.0000$\bullet$ .$4342\textpm.0065$5004±.0000$\bullet$ .3966±.0091$\bullet$ .4239±.0045$\bullet$ .4336±.0135$\bullet$ OOM .5408±.0016$\bullet$ .5649±.0476 OOM .5679±.0070
Precision .5688±.0000$\bullet$ .$5737\textpm.0037$5378±.0000$\bullet$ .5092±.0079$\bullet$ .5726±.0106$\bullet$ .5249±.0050$\bullet$ OOM .6711±.0042$\bullet$ .6524±.0297$\bullet$ OOM .7132±.0116
Recall .3782±.0000$\bullet$ .$3495\textpm.0094$4678±.0000$\bullet$ .3249±.0090$\bullet$ .3365±.0040$\bullet$ .3695±.0166$\bullet$ OOM .4528±.0008$\bullet$ .4995±.0557 OOM .4719±.0055
Purity .6961±.0000$\bullet$ .$6930\textpm.0095$6387±.0000$\bullet$ .5944±.0094$\bullet$ .6593±.0095$\bullet$ .6619±.0000$\bullet$ OOM .7190±.0022$\bullet$ .7057±.0163$\bullet$ OOM .7731±.0075
Time (s) 113.7 $82.2$ 16.4 87.4 7.7 2.9 OOM 195.6 202 OOM 103.5
@@
Figure 2: Clustering maps on India Pines (a) ground truth, (b) K-means, (c) FCM, (d) SSC, (e) SuperPCA, (f) SSSC, (g) S5C, (h) GCOT, (i) SGLSC, (j) DGAE, (k) NCSC, (l) SPGCC.
@@
Figure 3: Clustering maps on Salinas (a) ground truth, (b) K-means, (c) FCM, (d) SSC, (e) SuperPCA, (f) SSSC, (g) S5C, (h) SGLSC, (i) DGAE, (j) NCSC, (k) SPGCC.
@@
Figure 4: Clustering maps on Pavia University (a) ground truth, (b) K-means, (c) FCM, (d) SSC, (e) SuperPCA, (f) SSSC, (g) S5C, (h) SGLSC, (i) DGAE, (j) SPGCC.
@@ IV-C Comparison with Other Methods
According to the quantitative metrics and clustering maps, it could be easily observed that the proposed SPGCC method outperforms other compared methods on three datasets significantly, especially on India Pines. For example, OA obtained by SPGCC has an improvement of 8.8%, 2.4%, and 3.2% over the second-best method on the three datasets, respectively. Moreover, all the nine clustering metrics of SPGCC exceed those of the compared methods on India Pines. Only NMI and Recall on Salinas and Recall on Pavia University do not achieve the highest scores. Results of the significance test further validate the superiority of SPGCC (i.e., all metrics on India Pines, five metrics on Salinas, and six metrics on Pavia University have statistical superiority to the suboptimal methods).
@@
Compared with the state-of-the-art graph learning methods (i.e., SGLSC and DGAE), SPGCC performs better on all three datasets, with similar or less time cost. SPGCC uses 3-D and 2-D hybrid CNNs at the pre-training stage to extract high-order spatial and spectral features of HSI pixels, which is conducive to subsequent clustering, while SGLSC performs clustering on raw HSI features. Moreover, SPGCC is based on contrastive clustering and has an advantage over the GAE-based method because of the clustering-oriented optimization target (i.e., pulling superpixel representations of the same class close, while pushing different clusters’ representations away). In this manner, SPGCC gains clustering-friendly representations with better intra-class similarity and inter-class dissimilarity.
@@ IV-D Parameters Analysis
Our method mainly contains four hyperparameters: number of superpixels ($M$), learning rate ($\eta$), weight of clustering-center-level contrast loss ($\alpha$), and ratio of high-confidence samples ($\lambda$). They have been tuned for different datasets. The adopted values are shown in Table III.
@@
In Fig. 6, we show the effect of different $M$ on the running time of our algorithm. Settings of $M$ keep consistency with Fig. 5. Specifically, we fix the training epochs to 200 and record the running time. Due to the efficient parallelism of GCN, the increase in time consumption is lower than that of $M$.
The impact of $M$ on superpixel segmentation accuracy $SP_{acc}$ is also demonstrated in Fig. 6, which is defined as:
$SP_{acc}=\frac{\sum_{i=1}^{M}n_{i}^{dominant}}{\sum_{i=1}^{M}n_{i}^{labeled}},$ (21)
where $n_{i}^{dominant}$ and $n_{i}^{labeled}$ are the number of labeled pixels in the dominant class and total labeled pixels within the $i$-th superpixel, respectively. With the increase of $M$, $SP_{acc}$ shows a gradual upward trend, indicating that pixels within one superpixel are more likely to belong to the same class in a higher granularity partition. Since $SP_{acc}$ typically exceeds 99%, different $M$ have minimal impact on the accuracy of superpixel segmentation.
@@
2) Impact of $\lambda$: $\lambda$ controls the ratio of selected samples when recomputing high-confidence clustering centers. Fig. 7 shows the effect of $\lambda$ on clustering metrics. According to Fig. 7, the proposed SPGCC is quite robust to parameter $\lambda$. Specifically, fluctuations of the clustering performance on the India Pines dataset are slightly greater than those on Salinas and Pavia University. One possible reason is that the noise intensity on India Pines is larger than that on other datasets and the number of divided superpixels is smaller, so a larger $\lambda$ makes clustering centers more susceptible to noise samples, while a smaller $\lambda$ means only a small part of samples being selected which cannot fully represent centers of certain classes. Tuning results show that the optimal values of $\lambda$ for the three datasets are respectively 0.75, 0.55, and 0.25.
@@
TABLE VI: Ablation Study Results, OA is Repored.
@@
w/o pixel-level pre-training
.5354
.6783 .5049
w/o high-confidence centers
.6548
.8364 .6218
w/o pixel sampling augmentation
.6615
.8419 .6164
w/o model weight augmentation
.6416
.8247 .5954
w/o sample-level alignment
.6503
.8418 .5993
w/o clustering-center-level contrast
.6534
.8365 .6137
SPGCC .6759 .8447 .6268
TABLE VII: Different Data Augmentations for Superpixels, OA is Repored, Pixel Sampling Augmentation and Model Weight Augmentation are Abbreviated to “PSA” and “MWA”, Bold and Underline Denote the Optimal and Suboptimal Values Respectively.
Augmentation India Pines Salinas Pavia University
Node mask
.6274
.8225 .5801
Node noise
.6202
.8217 .5783
Node shuffling
.6401
.8230 .5822
Edge permutation
.6327
.8377 .5899
PSA
.6416
.8247 .5954
MWA .6615 .8419 .6164
PSA+MWA .6759 .8447 .6268
@@ IV-E Ablation Study
To validate the effectiveness of each module in the proposed SPGCC, including pixel-level pre-training, recomputed high-confidence clustering centers, two types of semantic-invariant superpixel augmentations (i.e., pixel sampling augmentation and model weight augmentation), two clustering-oriented training targets (i.e., sample-level alignment and clustering-center-level contrast), we conduct a series of ablation experiments. According to Table VI, the complete SPGCC achieves the best performance. In addition, we summarize the following four observations. First, pixel-level pre-training is very important and brings more than 10% performance improvement, indicating that high-order spectral and spatial features extracted from pre-training are more conducive to clustering. Second, recomputing clustering centers only with a certain proportion of high-confidence samples can alleviate the influence of noise samples and thus improve clustering performance. Third, using two types of semantic-invariant superpixel data augmentations simultaneously performs better than using only one, which is consistent with the current opinion in self-supervised contrastive learning that multiple reasonable data augmentations are beneficial to representation learning. Fourth, combining sample-level alignment with clustering-center-level contrast performs better, indicating that sample-level alignment helps to increase intra-class similarity while clustering-center-level contrast helps to increase inter-class dissimilarity, thus making representations easier to cluster.
@@ IV-F Different Superpixel-level Data Augmentations
We also compared the proposed two semantic-invariant superpixel augmentations, i.e., pixel sampling augmentation (PSA) and model weight augmentation (MWA), with general graph data augmentation methods[41] which may hurt the node semantic in Table VII. The compared data augmentations include node mask, node noise, node shuffling, and edge permutation. We tune the changing rate from 10% to 30% and report their best performance.
@@
IV-G Visualization of Superpixel Representations with t-SNE
$t$-distributed stochastic neighbor embedding (t-SNE)[53] is used for visualization of superpixel representations. We show the visualization results of the raw HSI features, pixel-level pre-trained features, and embeddings after SPGCC training in Fig. 9. The clustering performance of each type of feature is reported as the best result among ten times K-means.

### v2

@@ IV-B Experiment Setup
Dataset Metric (%) K-means FCMFCM SSC SuperPCA SSSC S5C GCOT SGLSC DGAE NCSC S3-ULDA SPGCC
[…]
Time (s) 22.6 $31.3$ 20.3 6.4 2.0 1.4 146.1 108.0 25.5 77.5 11.5 31.1
[…]
Time (s) 109.5 $155.2$ 11.7 92.1 12.4 2.4 OOM 47.0 360.2 2408.3 194.0 150.6
[…]
Time (s) 113.7 $82.2$ 16.4 87.4 7.7 2.9 OOM 195.6 202.0 OOM 374.5 103.5
@@
Figure 2: Clustering maps on India Pines (a) ground truth, (b) K-means, (c) FCM, (d) SSC, (e) SuperPCA, (f) SSSC, (g) S5C, (h) GCOT, (i) SGLSC, (j) DGAE, (k) NCSC, (l) S3-ULDA, (m) SPGCC.
@@
Figure 3: Clustering maps on Salinas (a) ground truth, (b) K-means, (c) FCM, (d) SSC, (e) SuperPCA, (f) SSSC, (g) S5C, (h) SGLSC, (i) DGAE, (j) NCSC, (k) S3-ULDA, (l) SPGCC.
@@
Figure 4: Clustering maps on Pavia University (a) ground truth, (b) K-means, (c) FCM, (d) SSC, (e) SuperPCA, (f) SSSC, (g) S5C, (h) SGLSC, (i) DGAE, (j) S3-ULDA, (k) SPGCC.
@@ IV-C Comparison with Other Methods
According to the quantitative metrics and clustering maps, it could be easily observed that the proposed SPGCC method outperforms other compared methods on three datasets significantly. For example, OA obtained by SPGCC has an improvement of 2.5%, 2.4%, and 2.7% over the second-best method on the three datasets, respectively. Results of the significance test further validate the superiority of SPGCC (i.e., eight metrics on India Pines, five metrics on Salinas, and four metrics on Pavia University have statistical superiority to the suboptimal values of different compared methods).
@@
Compared with the recent graph learning methods (i.e., SGLSC and DGAE), SPGCC performs better on all three datasets, with similar or less time cost. SPGCC uses 3-D and 2-D hybrid CNNs at the pre-training stage to extract high-order spatial and spectral features of HSI pixels, which is conducive to subsequent clustering, while SGLSC performs clustering on raw HSI features. Moreover, SPGCC is based on contrastive clustering and has an advantage over the GAE-based method because of the clustering-oriented optimization target (i.e., pulling superpixel representations of the same class close, while pushing different clusters’ representations away). In this manner, SPGCC gains clustering-friendly representations with better intra-class similarity and inter-class dissimilarity.
@@ IV-D Parameters Analysis
Our method mainly contains four hyperparameters: number of superpixels ($M$), weight of clustering-center-level contrast loss ($\alpha$), ratio of high-confidence samples ($\lambda$) and number of GCN layers ($L$). They have been tuned for different datasets. The adopted values are shown in Table III.
@@
Moreover, we validate the effect of $M$ on the running time and the superpixel segmentation accuracy in Fig. 6. Specifically, we fix the training epochs to 200 and record the running time. Due to the efficient parallelism of GCN, the increase in time consumption is lower than that of $M$. As for the superpixel segmentation accuracy, we define a metric called average superpixel segmentation accuracy $SP^{Avg}_{acc}$ as:
$SP^{Avg}_{acc}=\frac{\sum_{i=1}^{M}n_{i}^{dominant}}{\sum_{i=1}^{M}n_{i}^{labeled}},$ (21)
where $n_{i}^{dominant}$ and $n_{i}^{labeled}$ are the numbers of labeled pixels in the dominant class and total labeled pixels within the $i$-th superpixel, respectively. With the increase of $M$, $SP^{Avg}_{acc}$ shows a gradual upward trend, indicating that pixels within one superpixel are more likely to belong to the same class in a higher granularity partition. Since $SP^{Avg}_{acc}$ typically exceeds 99%, different $M$ have minimal impact on the accuracy of superpixel segmentation.
@@
Figure 9: The influence of different numbers of GCN layers $L$ on (a) India Pines, (b) Salinas, (c) Pavia University.
2) Impact of $\lambda$: $\lambda$ controls the ratio of selected samples when recomputing high-confidence clustering centers. According to Fig. 7, the proposed SPGCC is quite robust to parameter $\lambda$. Specifically, fluctuations of the clustering performance on the India Pines dataset are slightly greater than those on Salinas and Pavia University. One possible reason is that the noise intensity on India Pines is larger than that on other datasets and the number of divided superpixels is smaller, so a larger $\lambda$ makes clustering centers more susceptible to noise samples, while a smaller $\lambda$ means only a small part of samples being selected which cannot fully represent centers of certain classes. Tuning results show that the optimal values of $\lambda$ for the three datasets are respectively 0.75, 0.55, and 0.25.
@@
4) Impact of $L$: $L$ is the number of GCN layers. As shown in Fig. 9, it is evident that as $L$ increases, clustering metrics exhibit an increasing-then-decreasing trend. The best clustering performance is achieved when $L$ is set to 3 on all three datasets. Since one layer GCN can aggregate $1$-hop neighbors, small $L$ makes the model inaccessible to high-order neighbors, while large $L$ may lead to the over-smoothing problem, i.e., representations of different samples tend to be consistent and lack discriminability. Therefore, we empirically fix $L$ to 3.
TABLE VI: Ablation Study Results, OA (%) is Repored.
@@
w/o pixel-level pre-training 53.54 67.83 50.49
w/o high-confidence centers 65.48 83.64 62.18
w/o pixel sampling augmentation 66.15 84.19 61.64
w/o model weight augmentation 64.16 82.47 59.54
w/o sample-level alignment 65.03 84.18 59.93
w/o clustering-center-level contrast 65.34 83.65 61.37
SPGCC 67.59 84.47 62.68
TABLE VII: Different Data Augmentations for Superpixels, OA (%) is Repored, Pixel Sampling Augmentation and Model Weight Augmentation are Abbreviated to “PSA” and “MWA”, Bold and Underline Denote the Optimal and Suboptimal Values Respectively.
Type Augmentation India Pines Salinas Pavia University
Node-level Node mask 62.74 82.25 58.01
Node noise 62.02 82.17 57.83
Node shuffling 64.01 82.30 58.22
Edge-level Edge perturbation 63.27 83.77 58.99
Semantic-invariant PSA 64.16 82.47 59.54
MWA 66.15 84.19 61.64
PSA+MWA 67.59 84.47 62.68
Figure 10: OA (%) obtained by different contrastive learning methods with different temperature scaling parameters $\tau$ on (a) India Pines, (b) Salinas, (c) Pavia University.
@@ IV-E Ablation Study
To validate the effectiveness of each module in the proposed SPGCC, including pixel-level pre-training, recomputed high-confidence clustering centers, two types of semantic-invariant superpixel augmentations (i.e., pixel sampling augmentation and model weight augmentation), two clustering-oriented training targets (i.e., sample-level alignment and clustering-center-level contrast), we conduct a series of ablation experiments. According to Table VI, the complete SPGCC achieves the best performance. In addition, we summarize the following four observations. First, pixel-level pre-training is very important and brings more than 10% performance improvement, indicating that high-order spectral and spatial features extracted from pre-training are more conducive to clustering. Second, recomputing clustering centers only with a certain proportion of high-confidence samples can alleviate the influence of noise samples and thus improve clustering performance. Third, using two types of semantic-invariant superpixel data augmentations simultaneously performs better than only using one, which is consistent with the current opinion in self-supervised contrastive learning that multiple reasonable data augmentations are beneficial to representation learning. Fourth, combining sample-level alignment with clustering-center-level contrast performs better, indicating that sample-level alignment helps to increase intra-class similarity while clustering-center-level contrast helps to increase inter-class dissimilarity, thus making representations easier to cluster.
@@ IV-F Different Superpixel-level Data Augmentations
We also compared the proposed two semantic-invariant superpixel augmentations, i.e., pixel sampling augmentation (PSA) and model weight augmentation (MWA), with general graph data augmentation methods[48] which may hurt the node semantic in Table VII. The compared data augmentations include node mask, node noise, node shuffling, and edge perturbation. We tune the changing rate from 10% to 30% and report their best performance.
@@
IV-G Different Contrastive Learning Methods
Then, we compare the sample-level alignment (SLA) and clustering-center-level contrast (CLC) with other well-known contrastive learning methods, i.e., MOCO [33] and CC [35]. For fairness, we report the performance with different temperature scaling parameters $\tau$ for each method.
From Fig. 10, it can be observed that both MOCO and CC are inferior to the proposed SPGCC. This validates the shortcoming of the sample-level contrast strategy adopted in MOCO and CC, which may treat samples in the same class as negative samples, pushing them away in the representation space and disrupting intra-class similarity. While our method avoids this problem by combining SLA and CLC, leading to better clustering performance.
IV-H Visualization of Superpixel Representations with t-SNE
$t$-distributed stochastic neighbor embedding (t-SNE)[60] is used for visualization of superpixel representations. We show the visualization results of the raw HSI features, pixel-level pre-trained features, and embeddings after SPGCC training in Fig. 11.